OpenStudy (anonymous):

Matrix X-py = 4 -2x+y= q Find values of p and q for which system equations has a unique solution

OpenStudy (anonymous):

slopes of the two lines must not be equal!!

OpenStudy (anonymous):

\[{-1\over p}\ne{-2}\]

OpenStudy (anonymous):

"q" can be any real number

OpenStudy (anonymous):

x-py=4 -2x+y=q writing in the matrix form [■(1&-p@-2&1)][■(x@y)]=[■(4@q)] Or AX=B X=A-1 B |A|=|■(1&-p@-2&1)| |A|=1*1-(-p*-2) =1-2p For a unique solution |A| ≠0 i.e 1-2p≠0 1≠2p p≠1/(2 ), q can have any value